How many molecules of acetyl-CoA are produced by catabolism of the following fatty acids, and how many β oxidations are needed?

Part A

Lauric acid, CH3(CH2)10COOH

Express your answers as integers separated by a comma.

molecules of acetyl-CoA, β oxidations =

Part B

Decanoic acid, CH3(CH2)8COOH

Express your answers as integers separated by a comma.

molecules of acetyl-SCoA, β oxidations =
Answer #1

Lauric acid, CH3(CH2)10COOH, has 12 carbons.

Since the Acetyl group of the acetyl coA is formed by two carbons, we should divide the number of carbons in the acyl group between two.

Thus the number of molecules of acetyl-CoA = 12/2= 6

And the number of β oxidations is (n/2) -1 = 12/2-1=6-1=5

Decanoic acid, CH3(CH2)8COOH

Thus the number of molecules of acetyl-CoA = 10/2= 5

And the number of β oxidations is (n/2) -1 = 10/2-1=5-1=4